# On RoCK blocks of double covers of symmetric and alternating groups and the refined Broué conjecture

### v1

@@ 2 Preliminaries
The groups $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$. The double covers of $S_{n}$ (where $n$ is a non-negative integer) are given by
@@
We will use $\widetilde{S}_{n}$ to simultaneously represent both double covers. Denote by $\pi_{n}$ the surjective group homomorphism $\widetilde{S}_{n}\to S_{n}$, $z\mapsto 1$, $t_{i}\mapsto s_{i}$, where $s_{i}$ denotes the transposition $(i,i+1)\in S_{n}$. Then the double cover of $A_{n}$ is given by $\widetilde{A}_{n}:=\pi_{n}^{-1}(A_{n})$.
@@ 3 Descent and supermodules
3.2.
Let $R\subseteq R^{\prime}$ be two commutative rings. Let $A$ be an $R$-superalgebra and let $A^{\prime}:=R^{\prime}\otimes_{R}A$. Let $U^{\prime}$ be an $R^{\prime}$-supermodule. We said that $U^{\prime}$ is defined over $R$ or $U^{\prime}$ descends to $R$, if there is an $A$-supermodule $U$ such that $U^{\prime}\cong R^{\prime}\otimes_{R}U$ as $A^{\prime}$-supermodules. In this case, by the definition of an isomorphism of $R^{\prime}$-supermodules, we have $U^{\prime}_{\bar{0}}\cong R^{\prime}\otimes_{R}U_{\bar{0}}$ and $U^{\prime}_{\bar{1}}\cong R^{\prime}\otimes_{R}U_{\bar{1}}$ as $A^{\prime}_{\bar{0}}$-modules (recall from Proposition 3.1 that $A^{\prime}_{\bar{0}}=R^{\prime}\otimes_{R}A_{\bar{0}}$). The converse holds if $A$ has a super unit:
@@
Let $R\subseteq R^{\prime}$ be two commutative rings. Let $A$ be an $R$-superalgebra with a superunit, and let $A^{\prime}:=R^{\prime}\otimes_{R}A$. Let $U^{\prime}$ be an $R^{\prime}$-supermodule. If the $A^{\prime}_{\bar{0}}$-module $U^{\prime}_{\bar{0}}$ descends to $R$, then $U^{\prime}$ descends to $R$.
Proof. By assumption there is an $A_{\bar{0}}$-module $U_{0}$ such that $U^{\prime}_{\bar{0}}\cong R^{\prime}\otimes_{R}U_{\bar{0}}$ as $A^{\prime}_{\bar{0}}$-module. Let $U:=A\otimes_{A_{\bar{0}}}U_{\bar{0}}$. It is easy to check that $U$ is an $A$-supermodule. By [22, Lemma 3.7 (i)], we have $U^{\prime}\cong A^{\prime}\otimes_{A^{\prime}_{\bar{0}}}U^{\prime}_{\bar{0}}$. Hence we have
@@
Proof. By [23, Corollary 1.12.12], we have an $R^{\prime}$-module isomorphism
@@
Let $(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be a $p$-modular system which is an extension of $(K,\mathcal{O},k)$. Let $A$ and $B$ be $\mathcal{O}$-superalgebras, and let $A^{\prime}:=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}A$, $B^{\prime}:=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}B$. Let $M$ be an $(A,B)$-bisupermodule, NN a $(B,A)$-bisupermodule, $M^{\prime}:=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}M$ and $N^{\prime}\cong\mathcal{O}^{\prime}\otimes_{\mathcal{O}}N$. Then $M^{\prime}$ and $N^{\prime}$ induce a Morita superequivalence between $A^{\prime}$ and $B^{\prime}$ if and only if $M$ and $N$ induce a Morita superequivalence between $A$ and $B$. (Note that in this paper, all superspaces are defined to be finitely generated.) A similar statement holds with $\mathcal{O}^{\prime}$ and $\mathcal{O}$ replaced by $k^{\prime}$ and $k$ respectively.
Proof. One direction of the implications is trivial. Now assume that
$M^{\prime}\otimes_{B^{\prime}}N^{\prime}\cong A^{\prime}~~{\rm and}~~N^{\prime}\otimes_{A^{\prime}}M^{\prime}\cong B^{\prime}$ (3.1)
as $(A^{\prime},A^{\prime})$-bisupermodules and $(B^{\prime},B^{\prime})$-bisupermodules, respectively. Then by [19, Proposition 4.5],
$M\otimes_{B}N\cong A~~{\rm and}~~N\otimes_{A}M\cong B$ (3.2)
@@
Proposition 3.5 can be easily extend to the case for cross products (see 4.1) instead of superalgebras, and graded Morita equivalences (see [26, Definition 3.3]) or graded Rickard equivalences (see 4.1) instead of Morita superequivalences. The proof is almost the same.
3.7.
Projective supermodules and Heller translates. Let $R\in\{\mathcal{O},k\}$ and let $A$ be a superalgebra with a superunit. Following [22, §3.4], we call an $A$-supermodule $P$ a projective supermodule if $P$ is isomorphic to a direct summand of $A^{\oplus n}$ for some $n$. A projective cover of an $A$-supermodule is defined exactly like for modules. Our conditions on the superalgebra $A$ ensure that a projective cover of a finitely generated $A$-supermodule $M$ exists and is unique up to isomorphism. So we can define Heller translate $\Omega_{A}(M)$ of $M$ just like for modules.
@@
We refer to [23, Theorem 5.2.1] for the definition of Green correspondents. The following result is a descent criterion for Green correspondents.
Proposition 3.9.
Let $(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be a $p$-modular system which is an extension of $(K,\mathcal{O},k)$. Let $G$ be a finite group, $P$ a subgroup of $G$ and $H$ a subgroup of $G$ such that $N_{G}(P)\subseteq H$. Let $U^{\prime}$ be an $\mathcal{O}^{\prime}$-free indecomposable $\mathcal{O}^{\prime}G$-module having $P$ as a vertex, and let $V^{\prime}$ be an $\mathcal{O}^{\prime}H$-module which is the Green correspondent of $U^{\prime}$ with respect to $P$. Assume that $k$ is finite. Then $U^{\prime}$ descends to $\mathcal{O}$ if an only if $V^{\prime}$ descends to $\mathcal{O}$. A similar statement holds for $k^{\prime}$ and $k$ instead of $\mathcal{O}^{\prime}$ and $\mathcal{O}$ respectively.
Proof. Assume that $U^{\prime}$ descends to $\mathcal{O}$, that is, there exists an $\mathcal{O}G$-module $U$ such that $U^{\prime}\cong\mathcal{O}^{\prime}\otimes_{\mathcal{O}}U$. By definition, $V^{\prime}$ is the unique (up to isomorphism) direct summand of ${\rm Res}_{H}^{G}(U^{\prime})$ with vertex $P$. By [19, Lemma 6.1], we deduce that there is an indecomposable direct summand of $V$ of ${\rm Res}_{P}^{G}(U)$ such that $V^{\prime}\cong\mathcal{O}^{\prime}\otimes_{\mathcal{O}}V$.
@@
For example $G$ and $G_{\bar{0}}$ can be chosen to be $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$ ($n\geq 2$) respectively. For any subgroup $H$ of $G$ we will use $H_{\bar{0}}$ to denote $H\cap G_{\bar{0}}$ and $RH$ will inherit its superalgebra structure from that of $RG$.
@@
In [22, Theorem 3.39], Kleshchev and Livesey introduced a super version of the usual Green correspondence:
Theorem 3.11 ([22, Theorem 3.39]).
@@
Let $(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be a $p$-modular system which is an extension of $(K,\mathcal{O},k)$. Let $G$ be a finite group, $P$ a subgroup of $G$, and $H$ a subgroup of $G$ such that $N_{G}(P)\subseteq H$ and $H\nsubseteq G_{\bar{0}}$. Let $U^{\prime}$ be an $\mathcal{O}^{\prime}$-free module-indecomposable $\mathcal{O}^{\prime}G$-supermodule having $P$ as a vertex, and let $V^{\prime}$ be an $\mathcal{O}^{\prime}H$-module which is the super Green correspondent of $U^{\prime}$. Assume that $k$ is finite. Then $U$ descends to $\mathcal{O}$ if an only if $V$ descends to $\mathcal{O}$. A similar statement holds for $k^{\prime}$ and $k$ instead of $\mathcal{O}^{\prime}$ and $\mathcal{O}$ respectively.

### v5

@@ 2 Preliminaries
The groups $\widetilde{S}^{\eta}_{n}$ and $\widetilde{A}_{n}$. The double covers of $S_{n}$ (where $n$ is a non-negative integer) are given by
@@
We will use $\widetilde{S}^{\eta}_{n}$ ($\eta\in\{\pm\}$) to simultaneously represent both double covers. Denote by $\pi_{n}$ the surjective group homomorphism $\widetilde{S}^{\eta}_{n}\to S_{n}$, $z\mapsto 1$, $t_{i}\mapsto s_{i}$, where $s_{i}$ denotes the transposition $(i,i+1)\in S_{n}$. Then the double cover of $A_{n}$ is given by $\widetilde{A}_{n}:=\pi_{n}^{-1}(A_{n})$.
3 Descent of twisted wreath superproducts
3.1.
Let $R\in\{\mathcal{O},k\}$ and let $n$ be a positive integer. For $\eta\in\{\pm\}$, we define the $\eta$-twisted group superalgebra $\mathcal{T}^{\eta}_{n,R}$ of the symmetric group $S_{n}$ over $R$ to be the $R$-superalgebra given by odd generators $\tau_{1},\cdots,\tau_{n-1}$ subject to the relations
$\xi_{r}^{2}=\eta(-1)^{\frac{(p-1)}{2}},\quad\xi_{r}\xi_{s}=-\xi_{s}\xi_{r}\text{ if }|r-s|>1,\quad(\xi_{r}\xi_{r+1})^{3}=1.$
For each $w\in S_{n}$, choosing a reduced expression $w=s_{r_{1}}\cdots s_{r_{k}}$ in terms of simple transpositions, we define $\xi_{w}:=\xi_{r_{1}}\cdots\xi_{r_{k}}$ (which depends up to a sign on the choice of a reduced expression).
Proposition 3.2 (Descent of $\mathcal{T}_{n,R}$).
Let $(R,R_{p})\in\{(\mathcal{O},\mathbb{Z}_{p}),(k,\mathbb{F}_{p})\}$. If $\sqrt{-1}\in R$, then there is an isomorphism of $R$-superalgebras $\mathcal{T}^{\eta}_{n,R}\cong\mathcal{T}_{n,R}$ sending $\eta^{i}(-1)^{\frac{(p-1)i}{2}}\sqrt{\eta(-1)^{\frac{(p-1)}{2}}}\xi_{i}$ to $\tau_{i}$. Hence in this case, we have an isomorphism of $R$-superalgebras $R\otimes_{R_{p}}\mathcal{T}^{\eta}_{n,R_{p}}\cong\mathcal{T}_{n,R}$ sending $\eta^{i}(-1)^{\frac{(p-1)i}{2}}\sqrt{\eta(-1)^{\frac{(p-1)}{2}}}\otimes\xi_{i}$ to $\tau_{i}$. In particular, $\{\xi_{w}\mid w\in S_{n}\}$ is an $R_{p}$-basis of $\mathcal{T}_{n,R_{p}}^{\eta}$.
[…]
Definition 3.3.
Given an $R$-superalgebra AA, we define the $\eta$-twisted wreath superproduct $A\wr_{s}\mathcal{T}^{\eta}_{n,R}$ as the free product $A^{\otimes n}\star\mathcal{T}^{\eta}_{n,R}$ of superalgebras subject to the relations
$\xi_{r}\,(a_{1}\otimes\dots\otimes a_{n})=(-1)^{\sum_{u\neq r,r+1}|a_{u}|}\left({}^{s_{r}}(a_{1}\otimes\dots\otimes a_{n})\,\xi_{r}\right)$
Proposition 3.4.
Let $(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be a $p$-modular system which is an extension of $(K,\mathcal{O},k)$ such that $\sqrt{-1}\in\mathcal{O}^{\prime}$. Let $(R^{\prime},R)\in\{(\mathcal{O}^{\prime},\mathcal{O}),(k^{\prime},k)\}$, $n$ a positive integer, and $A^{\prime}$ an $R^{\prime}$-superalgebra. Then the following hold:
  1. 1.
    The isomorphism $\mathcal{T}^{\eta}_{n,R^{\prime}}\cong\mathcal{T}_{n,R^{\prime}}$ in Proposition 3.2 extends to an isomorphism of $R^{\prime}$-superalgebras $A^{\prime}\wr_{s}\mathcal{T}^{\eta}_{n,R^{\prime}}\cong A^{\prime}\wr_{s}\mathcal{T}_{n,R^{\prime}}$.
  2. 2.
    If $A^{\prime}\cong R^{\prime}\otimes_{R}A$ for some $R$-superalgebra AA, then $R^{\prime}\otimes_{R}(A\wr_{s}\mathcal{T}^{\eta}_{n,R})\cong A^{\prime}\wr_{s}\mathcal{T}_{n,R^{\prime}}$.
  3. 3.
    We have a decomposition
    $A\wr_{s}\mathcal{T}^{\eta}_{n,R}=\bigoplus_{w\in S_{n}}A^{\otimes n}\xi_{w}.$ (3.1)
    Hence $A\wr_{s}\mathcal{T}^{\eta}_{n,R}$ is an $S_{n}$-graded crossed $R$-superproduct.
Proof. The $\eta^{i}(-1)^{\frac{(p-1)i}{2}}\sqrt{\eta(-1)^{\frac{(p-1)}{2}}}\xi_{i}$-conjugation action and the $\tau_{i}$-conjugation action on an element $(a_{1}\otimes\dots\otimes a_{n})\in A^{\prime\otimes n}$ are the same, hence statement (i) holds. By statement (i) we have
$R^{\prime}\otimes_{R}(A\wr_{s}\mathcal{T}^{\eta}_{n,R})\cong(R^{\prime}\otimes_{R}A)\wr_{s}(R^{\prime}\otimes_{R}\mathcal{T}^{\eta}_{n,R})\cong A^{\prime}\wr_{s}\mathcal{T}^{\eta}_{n,R^{\prime}}\cong A^{\prime}\wr_{s}\mathcal{T}_{n,R^{\prime}}$
as $R^{\prime}$-superalgebras, whence (ii). By Proposition 2.8, we have
$A^{\prime}\wr_{s}\mathcal{T}_{n,R^{\prime}}=\bigoplus_{w\in S_{n}}A^{\prime\otimes n}\tau_{w}.$
Since $\eta^{i}(-1)^{\frac{(p-1)i}{2}}\sqrt{\eta(-1)^{\frac{(p-1)}{2}}}$ is invertible in $R^{\prime}$, by statement (ii), we have
$A^{\prime}\wr_{s}\mathcal{T}^{\eta}_{n,R^{\prime}}=\bigoplus_{w\in S_{n}}A^{\prime\otimes n}\xi_{w}.$
[…]
3.5.
Keep the notation of Proposition 3.4. Given the $S_{n}$-grading in (3.1), recall the definition of $(A\wr_{s}\mathcal{T}^{\eta}_{n,R},A\wr_{s}\mathcal{T}^{\eta}_{n,R})_{S_{n}}$ from (2.2). Let $M$ be an $(A,A)$-bisupermodule. Analogous to (2.5), we denote by $M^{\boxtimes n}_{S_{n}}$ the $(A\wr_{s}\mathcal{T}^{\eta}_{n,R},A\wr_{s}\mathcal{T}^{\eta}_{n,R})_{S_{n}}$-supermodule given by extending the $(A^{\otimes n},A^{\otimes n})$-bisupermodule $M^{\boxtimes n}:=\underbrace{M\boxtimes_{R}\cdots\boxtimes_{R}M}_{n}$, via
[…]
for all $m_{1},\cdots,m_{n}\in M$ and $1\leq r\leq n-1$.
Analogous to (2.6), we now define
$M\wr_{s}\mathcal{T}^{\eta}_{n,R}:={\rm Ind}_{(A\wr_{s}\mathcal{T}^{\eta}_{n,R},A\wr_{s}\mathcal{T}^{\eta}_{n,R})_{S_{n}}}^{(A\wr_{s}\mathcal{T}^{\eta}_{n,R})\otimes_{R}(A\wr_{s}\mathcal{T}^{\eta}_{n,R})^{\rm sop}}(M^{\boxtimes n}_{S_{n}}).$ (3.3)
Proposition 3.6.
Let $(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be a $p$-modular system which is an extension of $(K,\mathcal{O},k)$ such that $\sqrt{-1}\in\mathcal{O}^{\prime}$. Let $(R^{\prime},R)\in\{(\mathcal{O}^{\prime},\mathcal{O}),(k^{\prime},k)\}$, $n$ a positive integer, $A^{\prime}$ an $R^{\prime}$-superalgebra and $M^{\prime}$ an $(A^{\prime},A^{\prime})$-bisupermodule. Through the isomorphism $A^{\prime}\wr_{s}\mathcal{T}^{\eta}_{n,R^{\prime}}\cong A^{\prime}\wr_{s}\mathcal{T}_{n,R^{\prime}}$, we regard the $(A^{\prime}\wr_{s}\mathcal{T}_{n,R^{\prime}},A^{\prime}\wr_{s}\mathcal{T}_{n,R^{\prime}})$-bisupermodule $M^{\prime}\wr_{s}\mathcal{T}_{n,R^{\prime}}$ as an $(A^{\prime}\wr_{s}\mathcal{T}^{\eta}_{n,R^{\prime}},A^{\prime}\wr_{s}\mathcal{T}^{\eta}_{n,R^{\prime}})$-bisupermodule. Then the following hold:
  1. 1.
    We have an isomorphism of $(A^{\prime}\wr_{s}\mathcal{T}^{\eta}_{n,R^{\prime}},A^{\prime}\wr_{s}\mathcal{T}^{\eta}_{n,R^{\prime}})$-bisupermodules $M^{\prime}\wr_{s}\mathcal{T}^{\eta}_{n,R^{\prime}}\cong M^{\prime}\wr_{s}\mathcal{T}_{n,R^{\prime}}$.
  2. 2.
    If $A^{\prime}\cong R^{\prime}\otimes_{R}A$ for some $R$-superalgebra AA and $M^{\prime}\cong R^{\prime}\otimes_{R}M$ for some $(A,A)$-bisupermodule $M$, then $R^{\prime}\otimes_{R}(M\wr_{s}\mathcal{T}^{\eta}_{n,R})\cong M^{\prime}\wr_{s}\mathcal{T}_{n,R^{\prime}}$ as $(A^{\prime}\wr_{s}\mathcal{T}^{\eta}_{n,R^{\prime}},A^{\prime}\wr_{s}\mathcal{T}^{\eta}_{n,R^{\prime}})$-bisupermodules.
[…]
3.7.
Let $R\in\{\mathcal{O},k\}$, $n$ a positive integer and $A$ an $R$-superalgebra with a superunit $u$. If $C_{2}$ is a cyclic group of order $2$ with generator $g$, we can give $A$ the structure of a $C_{2}$-graded crossed $R$-superproduct with graded components $A_{g^{\varepsilon}}=A_{\bar{0}}u^{\varepsilon}=A_{\varepsilon}$ for $\varepsilon\in\mathbb{Z}/2\mathbb{Z}$. For any $n\in\mathbb{Z}_{>0}$, again we consider the wreath product
where $g_{r}$ is the generator of the $r$-th $C_{2}$ factor in the base group $C_{2}^{\times n}$. Then we can give $A\wr_{s}\mathcal{T}^{\eta}_{n,R}$ the structure of a crossed product of $A_{\bar{0}}^{\otimes n}$ and $C_{2}\wr S_{n}$ (see 5.1) with graded components
$(A\wr_{s}\mathcal{T}^{\eta}_{n,R})_{g_{1}^{\varepsilon_{1}}\cdots g_{n}^{\varepsilon_{n}}w}=A_{\bar{0}}^{\otimes n}u_{1}^{\varepsilon_{1}}\dots u_{n}^{\varepsilon_{n}}\tau_{w}$
where each $u_{r}$ is $1^{\otimes(r-1)}\otimes u\otimes 1^{\otimes(n-r)}$.
$(A\wr_{s}\mathcal{T}^{\eta}_{n,R},B\wr_{s}\mathcal{T}^{\eta}_{n,R})_{C_{2}\wr S_{n}}\subseteq(A\wr_{s}\mathcal{T}^{\eta}_{n,R})\otimes_{R}(B\wr_{s}\mathcal{T}^{\eta}_{n,R})^{\rm sop}.$
Setting $U_{r}:=(u_{A})_{r}\otimes(u_{B}^{-1})_{r}$, for $1\leq r\leq n$, we have that $A_{\bar{0}}^{\otimes n}\otimes_{R}(B_{\bar{0}}^{\otimes n})^{\rm sop}$ together with the $U_{r}$’s generate a subsuperalgebra of $(A\wr_{s}\mathcal{T}^{\eta}_{n,R},B\wr_{s}\mathcal{T}^{\eta}_{n,R})_{C_{2}\wr_{s}S_{n}}$ isomorphic to $((A,B)_{C_{2}})^{\otimes n}$. We identify the two superalgebras. Then $(A\wr_{s}\mathcal{T}^{\eta}_{n,R},B\wr_{s}\mathcal{T}^{\eta}_{n,R})_{C_{2}\wr_{s}S_{n}}$ is generated by $((A,B)_{C_{2}})^{\otimes n}$ and $\{\Xi_{r}:=\xi_{r}\otimes\xi_{r}^{-1}\mid 1\leq r\leq n-1\}$ subject to the following relations:
[…]
where $\underline{a}=a_{1}\otimes\dots\otimes a_{n}\in A_{\bar{0}}^{\otimes n}$ and $\underline{b}=b_{1}\otimes\dots\otimes b_{n}\in(B_{\bar{0}}^{\otimes n})^{\rm op}$.
Proposition 3.8 (a slight modification of [22, Proposition 5.1.14]).
Let $R\in\{\mathcal{O},k\}$, let $n$ be a positive integer, and let $A$ and $B$ be $R$-superalgebras with superunit. If $X$ is a complex of $A_{\bar{0}}$-$B_{\bar{0}}$-bimodules that extends to a complex of $(A,B)_{C_{2}}$-modules, then $X^{\otimes n}$ extends to a complex of $(A\wr_{s}\mathcal{T}^{\eta}_{n,R},B\wr_{s}\mathcal{T}^{\eta}_{n,R})_{C_{2}\wr S_{n}}$-modules.
[…]
@@ 4 Descent and supermodules
4.2.
Let $R\subseteq R^{\prime}$ be two commutative rings. Let $A$ be an $R$-superalgebra and let $A^{\prime}:=R^{\prime}\otimes_{R}A$. Let $U^{\prime}$ be an $R^{\prime}$-supermodule. We say that $U^{\prime}$ is defined over $R$ or $U^{\prime}$ descends to $R$, if there is an $A$-supermodule $U$ such that $U^{\prime}\cong R^{\prime}\otimes_{R}U$ as $A^{\prime}$-supermodules. In this case, by the definition of an isomorphism of $R^{\prime}$-supermodules, we have $U^{\prime}_{\bar{0}}\cong R^{\prime}\otimes_{R}U_{\bar{0}}$ and $U^{\prime}_{\bar{1}}\cong R^{\prime}\otimes_{R}U_{\bar{1}}$ as $A^{\prime}_{\bar{0}}$-modules (recall from Proposition 4.1 that $A^{\prime}_{\bar{0}}=R^{\prime}\otimes_{R}A_{\bar{0}}$). The converse holds if $A$ has a super unit:
@@
Let $R\subseteq R^{\prime}$ be two commutative rings. Let $A$ be an $R$-superalgebra with a superunit, and let $A^{\prime}:=R^{\prime}\otimes_{R}A$. Let $U^{\prime}$ be an $A^{\prime}$-supermodule. If the $A^{\prime}_{\bar{0}}$-module $U^{\prime}_{\bar{0}}$ descends to $R$, then $U^{\prime}$ descends to $R$.
Proof. By assumption, there is an $A_{\bar{0}}$-module $U_{0}$ such that $U^{\prime}_{\bar{0}}\cong R^{\prime}\otimes_{R}U_{\bar{0}}$ as $A^{\prime}_{\bar{0}}$-module. Let $U:=A\otimes_{A_{\bar{0}}}U_{\bar{0}}$. It is easy to check that $U$ is an $A$-supermodule. By [23, Lemma 3.7 (i)], we have $U^{\prime}\cong A^{\prime}\otimes_{A^{\prime}_{\bar{0}}}U^{\prime}_{\bar{0}}$. Hence we have
@@
Proof. By [24, Corollary 1.12.12], we have an $R^{\prime}$-module isomorphism
@@
Let $(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be a $p$-modular system which is an extension of $(K,\mathcal{O},k)$. Let $A$ and $B$ be symmetric $\mathcal{O}$-superalgebras, and let $A^{\prime}:=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}A$, $B^{\prime}:=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}B$. Let $M$ be an $(A,B)$-bisupermodule, and let $M^{\prime}:=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}M$. Then $M^{\prime}$ and $M^{\prime*}$ induce a Morita superequivalence between $A^{\prime}$ and $B^{\prime}$ if and only if $M$ and $M^{*}$ induce a Morita superequivalence between $A$ and $B$. (Note that in this paper, all superspaces are defined to be finitely generated.) A similar statement holds with $\mathcal{O}^{\prime}$ and $\mathcal{O}$ replaced by $k^{\prime}$ and $k$ respectively.
Proof. One direction of the implications is trivial. Now assume that we have an $(A^{\prime},A^{\prime})$-bisupermodule isomorphism $f^{\prime}:M^{\prime}\otimes_{B^{\prime}}M^{\prime*}\cong A^{\prime}$ and a $(B^{\prime},B^{\prime})$-bisupermodule isomorphism $g^{\prime}:M^{\prime*}\otimes_{A^{\prime}}M^{\prime}\cong B^{\prime}$. Then by [20, Proposition 4.5], we have an $(A,A)$-bimodule isomorphism $f:M\otimes_{B}M^{*}\cong A$ and a $(B,B)$-bimodule isomorphism $g:M^{*}\otimes_{A}M\cong B$. Moreover, by the proof of [20, Proposition 4.5], we have $f^{\prime}={\rm id}_{\mathcal{O}^{\prime}}\otimes f$ as $(A^{\prime},A^{\prime})$-bimodule isomorphisms, and $g^{\prime}={\rm id}_{\mathcal{O}^{\prime}}\otimes g$ as $(B^{\prime},B^{\prime})$-bimodule isomorphisms. By Propositions 4.1 and 4.4, we have
$A^{\prime}_{\bar{0}}=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}A_{\bar{0}},~~A^{\prime}_{\bar{1}}=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}A_{\bar{1}},$
$(M^{\prime}\otimes_{B^{\prime}}M^{\prime*})_{\bar{0}}=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}(M\otimes_{B}M^{\prime*})_{\bar{0}}$
and
$(M^{\prime}\otimes_{B^{\prime}}M^{\prime*})_{\bar{1}}=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}(M\otimes_{B}M^{*})_{\bar{1}}.$
@@
Proposition 4.5 can easily be extended to the case of crossed products (see 5.1) instead of superalgebras, and to graded Morita equivalences (see [26, Definition 3.3]) or graded Rickard equivalences (see 5.1) instead of Morita superequivalences. The proof is almost the same.
4.7.
Projective supermodules and Heller translates. Let $R\in\{\mathcal{O},k\}$ and let $A$ be a superalgebra with a superunit. Following [23, §3.4], we call an $A$-supermodule $P$ a projective supermodule if $P$ is isomorphic to a direct summand of $A^{\oplus n}$ for some $n$. A projective cover of an $A$-supermodule is defined exactly like for modules. Our conditions on the superalgebra $A$ ensure that a projective cover of a finitely generated $A$-supermodule $M$ exists and is unique up to isomorphism. So we can define Heller translate $\Omega_{A}(M)$ of $M$ just like for modules.
@@
We refer to [24, Theorem 5.2.1] for the definition of Green correspondents. The following is a descent criterion for Green correspondents.
Proposition 4.9 (Descent of Green correspondents).
Let $(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be a $p$-modular system which is an extension of $(K,\mathcal{O},k)$. Let $G$ be a finite group, $P$ a subgroup of $G$ and $H$ a subgroup of $G$ such that $N_{G}(P)\subseteq H$. Let $U^{\prime}$ be an $\mathcal{O}^{\prime}$-free indecomposable $\mathcal{O}^{\prime}G$-module having $P$ as a vertex, and let $V^{\prime}$ be an $\mathcal{O}^{\prime}H$-module which is the Green correspondent of $U^{\prime}$ with respect to $P$. Assume that $\mathcal{O}^{\prime}$ is finitely generated as $\mathcal{O}$-module. Then $U^{\prime}$ descends to $\mathcal{O}$ if an only if $V^{\prime}$ descends to $\mathcal{O}$. A similar statement holds for $k^{\prime}$ and $k$ instead of $\mathcal{O}^{\prime}$ and $\mathcal{O}$ respectively.
Proof. Assume that $U^{\prime}$ descends to $\mathcal{O}$, that is, there exists an $\mathcal{O}G$-module $U$ such that $U^{\prime}\cong\mathcal{O}^{\prime}\otimes_{\mathcal{O}}U$. By definition, $V^{\prime}$ is up to isomorphism the unique direct summand of ${\rm Res}_{H}^{G}(U^{\prime})$ with vertex $P$. By [20, Lemma 5.1], we deduce that there is an indecomposable direct summand $V$ of ${\rm Res}_{P}^{G}(U)$ such that $V^{\prime}\cong\mathcal{O}^{\prime}\otimes_{\mathcal{O}}V$.
@@
For example $G$ and $G_{\bar{0}}$ can be chosen to be $\widetilde{S}^{\eta}_{n}$ and $\widetilde{A}_{n}$ ($n\geq 2$) respectively. For any subgroup $H$ of $G$ we will use $H_{\bar{0}}$ to denote $H\cap G_{\bar{0}}$ and $RH$ will inherit its superalgebra structure from that of $RG$.
@@
In [23, Theorem 3.39], Kleshchev and Livesey introduced a super version of the usual Green correspondence:
Theorem 4.11 ([23, Theorem 3.39]).
@@
Let $(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be a $p$-modular system which is an extension of $(K,\mathcal{O},k)$. Let $G$ be a finite group, $P$ a subgroup of $G$, and $H$ a subgroup of $G$ such that $N_{G}(P)\subseteq H$ and $H\nsubseteq G_{\bar{0}}$. Let $U^{\prime}$ be an $\mathcal{O}^{\prime}$-free module-indecomposable $\mathcal{O}^{\prime}G$-supermodule having $P$ as a vertex, and let $V^{\prime}$ be an $\mathcal{O}^{\prime}H$-module which is the super Green correspondent of $U^{\prime}$. Assume that $\mathcal{O}^{\prime}$ is finitely generated as $\mathcal{O}$-module. Then $U$ descends to $\mathcal{O}$ if an only if $V$ descends to $\mathcal{O}$. A similar statement holds for $k^{\prime}$ and $k$ instead of $\mathcal{O}^{\prime}$ and $\mathcal{O}$ respectively.